# Rainbow PO: A Unified Framework for Combining Improvements in Preference Optimization

### v1

@@ 5 Conclusion
In this paper, we propose RainbowPO, a comprehensive framework that demystifies and enhances existing DPO methods through the integration of key components into a unified objective. Our findings highlight the effectiveness of length normalization, reference policy mixing, and contextual scaling, while also highlighting the promise of warm-up adjustments. However, the selective application of rejection sampling and home advantage is not providing incremental improvements when paired with the other methods. By demonstrating that these enhancements can coexist within a single algorithm to achieve state-of-the-art performance, we pave the way for future research and practical applications. We aim for this work to serve as a foundation for refining DPO methodologies and to inspire further exploration of untested components for integrated agents.
Acknowledgements
We thank Stephen Rawls, Supriyo Chakraborty and Kartik Balasubramaniam for helpful feedbacks and pointers to relevant works. We also thank Benjamin Therien, Ashwinee Panda, Zain Sarwar, Akshaj Kumar Veldanda and Andrei Mircea for helpful discussions.
Wenpin Tang and Hanyang Zhao are supported by NSF grant DMS-2206038, a start-up grant at Columbia University, and the Columbia Innovation Hub grant. The works of Hanyang Zhao and David D. Yao are part of a Columbia-CityU/HK collaborative project that is supported by InnotHK Initiative, The Government of the HKSAR and the AIFT Lab.
@@ A.1 RLHF
$\max_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\left[\mathbb{E}_{y\sim\pi(y\mid x)}\left[r_{\psi_{*}}(x,y)\right]-\beta\mathrm{KL}\left(\pi(\cdot\mid x)\|\pi_{\mathrm{ref}}(\cdot\mid x)\right)\right]$ (17)
@@ A.2 DPO
Appendix B Proofs and Details
B.1 Proof of ORPO upper bound in equation 4
Theorem 1.
$\Delta_{\theta}(x,y^{w},y^{l})=\frac{1}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{1}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)\geq 0$
$\mathcal{L}_{\text{ORPO-PO}}\left(\pi_{\theta}\right)=-\mathbb{E}_{(x,y^{w},y^{l})\sim\mathcal{D}}\left[\log\sigma\left(\log\frac{p_{\theta}(y_{w}|x)}{1-p_{\theta}(y_{w}|x)}-\log\frac{p_{\theta}(y_{l}|x)}{1-p_{\theta}(y_{l}|x)}\right)\right],$ (20)
has an upper bound such that
Proof.
Taking gradient of $f(x)$ with respect to $x$, we have:
B.2 Mathematical explanation of Different XPOs
Most xPOs’ categories are straightforward by directly checking their loss objective, see Table 7. Less explainable is ORPO, which we have proved in Theorem 1.
Appendix C Miscellaneous
Method Objective
DPO $-\log\sigma\left(\beta_{\text{reg}}\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta_{\text{reg}}\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$
IPO $\left(\beta_{\text{reg}}\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta_{\text{reg}}\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\frac{1}{2}\right)^{2}$
$f$-DPO $-\log\sigma\left(\beta_{\text{reg}}f^{\prime}\left(\frac{\pi_{\bm{\theta}}\left(y_{w}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{w}\mid x\right)}\right)-\beta_{\text{reg}}f^{\prime}\left(\frac{\pi_{\bm{\theta}}\left(y_{l}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{l}\mid x\right)}\right)\right)$
KTO $-\lambda_{w}\sigma\left(\beta_{\text{reg}}\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-z_{\text{ref}}\right)-\lambda_{l}\sigma\left(z_{\text{ref}}-\beta_{\text{reg}}\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right),\,$
$\text{where}\,\,z_{\text{ref}}=\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\beta_{\text{reg}}\text{KL}\left(\pi_{\theta}(y|x)||\pi_{\text{ref}}(y|x)\right)\right]$
ODPO $-\log\sigma\left(\beta_{\text{reg}}\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta_{\text{reg}}\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\Delta_{r}(x)\right)$
Mallows-DPO $-\log\sigma\left(\phi(x)\left[\beta_{\text{reg}}\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta_{\text{reg}}\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right]\right)$
R-DPO $-\log\sigma\left(\beta_{\text{reg}}\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta_{\text{reg}}\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\left(\alpha|y_{w}|-\alpha|y_{l}|\right)\right)$
CPO $-\log p_{\theta}(y_{w}|x)-\log\sigma\left(\beta_{\text{reg}}\log\pi_{\theta}(y_{w}|x)-\beta_{\text{reg}}\log\pi_{\theta}(y_{l}|x)\right)$
SimPO $-\log\sigma\left(\frac{\beta_{\text{reg}}}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta_{\text{reg}}}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)$
@@ Appendix C Miscellaneous
Preference Dataset $\mathcal{D}_{\text{pref}}$ Dataset with a set of preferred and dispreferred.
responses to train a reward model.
SFT Dataset $\mathcal{D}_{\text{sft}}$ Dataset with a set of input and label for supervised
fine-tuning.
@@
Regularization Hyper-parameters $\alpha,\beta$ Regularization Hyper-parameters for preference tuning.
@@
Appendix D Experimental details

### v2

@@ 5 Conclusion
In this paper, we propose RainbowPO, a comprehensive framework that demystifies and enhances existing DPO methods through the integration of key components into a unified objective. Our findings highlight the effectiveness of length normalization, reference policy mixing, and contextual scaling. However, the selective application of rejection sampling and home advantage is not providing incremental improvements either individually or when paired with the other methods. By demonstrating that these enhancements can coexist within a single algorithm to achieve state-of-the-art performance for fine-tuning SOTA commercial LLM, we pave the way for future research and practical applications. We aim for this work to serve as a foundation for refining DPO methodologies and to inspire further exploration of untested components for integrated agents.
Acknowledgments
We thank Stephen Rawls, Supriyo Chakraborty and Kartik Balasubramaniam for helpful feedbacks and pointers to relevant works. We also thank Benjamin Therien, Ashwinee Panda, Zain Sarwar, Akshaj Kumar Veldanda, and Andrei Mircea for helpful discussions.
Wenpin Tang and Hanyang Zhao are supported by NSF grant DMS-2206038, a start-up grant at Columbia University, and the Columbia Innovation Hub grant. The works of Hanyang Zhao and David D. Yao are part of a Columbia-CityU/HK collaborative project that is supported by InnoHK Initiative, The Government of the HKSAR and the AIFT Lab.
@@ A.1 RLHF
$\max_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\left[\mathbb{E}_{y\sim\pi(y\mid x)}\left[r_{\psi_{*}}(x,y)\right]-\beta\mathrm{KL}\left(\pi(\cdot\mid x)\|\pi_{\mathrm{ref}}(\cdot\mid x)\right)\right],$ (20)
@@ A.2 DPO
Appendix B Miscellaneous
@@
Preference Dataset $\mathcal{D}$ Dataset with a set of preferred and dispreferred responses.
Preference Dataset by RSO $\mathcal{D}_{\text{RSO}}$ Dataset with a set of preferred and dispreferred responses
sampled by Rejection Sampling.
@@
Regularization Hyper-parameter $\beta$ Regularization Hyper-parameter for preference tuning.
Mixing Hyper-parameter $\alpha$ Our proposed mixing coefficient for better reference policy.
@@
Appendix C Mathematical explanation of Different XPOs
Here we first list popular xPOs variants in the literature as in  Meng et al. (2024) and Winata et al. (2024).
Method Objective
DPO $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$
IPO $\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\frac{1}{2}\right)^{2}$
$f$-DPO $-\log\sigma\left(\beta f^{\prime}\left(\frac{\pi_{\bm{\theta}}\left(y_{w}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{w}\mid x\right)}\right)-\beta f^{\prime}\left(\frac{\pi_{\bm{\theta}}\left(y_{l}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{l}\mid x\right)}\right)\right)$
KTO $-\lambda_{w}\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-z_{\text{ref}}\right)-\lambda_{l}\sigma\left(z_{\text{ref}}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right),\,$
$\text{where}\,\,z_{\text{ref}}=\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\beta\text{KL}\left(\pi_{\theta}(y|x)||\pi_{\text{ref}}(y|x)\right)\right]$
ODPO $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\Delta_{r}(x)\right)$
MallowsPO $-\log\sigma\left(\phi(x)\left[\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right]\right)$
R-DPO $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\left(\alpha|y_{w}|-\alpha|y_{l}|\right)\right)$
CPO $-\log p_{\theta}(y_{w}|x)-\log\sigma\left(\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)\right)$
SimPO $-\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)$
Next we include the categorization of different methods, which appeared earlier in Table 9, and explain detailedly the reasons after.
DPO $\times$ logistic $\times$ SFT $\times$ $\times$ $\times$
SLiC-HF $\times$ hinge $\times$ SFT $\times$ $\times$ $\checkmark$
IPO $\times$ square $\times$ SFT $\times$ $\times$ $\times$
CPO $\times$ logistic $\times$ Free $\times$ $\times$ $\checkmark$
RSO $\times$ logistic / hinge $\times$ SFT $\times$ $\checkmark$ $\times$
ODPO $\times$ logistic $\checkmark$ SFT $\times$ $\times$ $\times$
ORPO $\checkmark$ logistic $\times$ Free implicitly $\times$ $\checkmark$
WPO $\times$ logistic $\times$ SFT $\checkmark$ $\times$ $\times$
MallowsPO $\times$ logistic $\times$ SFT $\checkmark$ $\times$ $\times$
SimPO $\checkmark$ logistic $\checkmark$ Free $\times$ $\times$ $\times$
[…]
C.1 Direct Preference Optimization (DPO)
The loss of DPO (Rafailov et al., 2023) is:
$\mathcal{L}_{\mathrm{DPO}}\left(\pi_{\theta};\pi_{\mathrm{ref}}\right):=-\mathbb{E}_{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(y_{w}\mid x\right)}{\pi_{\text{ref }}\left(y_{w}\mid x\right)}-\beta\log\frac{\pi_{\theta}\left(y_{l}\mid x\right)}{\pi_{\text{ref }}\left(y_{l}\mid x\right)}\right)\right],$ (23)
which can be seen that, as the baseline methods we investigate in this paper, no length normalization is adopted, link function $-\log\sigma$ adopts the logistic function, home advantage can be seen as None or 0, reference policy is the SFT policy to be aligned, no contextual scaling, the dataset is the hybrid or offline dataset thus there is no rejection sampling. Also there is no SFT loss.
C.2 Sequence Likelihood Calibration from Human Feedback (SLiC-HF)
The loss of SliC-HF (Zhao et al., 2023) is ($y_{\text{ref}}$ refers to answer generated by $\pi_{\text{ref}}$):
$\mathcal{L}_{\text{SLiC}}{(\pi_{\theta};\pi_{\mathrm{ref}})}=\mathbb{E}_{\left(x,y_{w},y_{l}\right)}\underbrace{\max\left(0,\delta-\log\pi_{\theta}(y_{w}|x)+\log\pi_{\theta}(y_{l}|x)\right)}_{\text{rank calibration loss}}\underbrace{-\lambda\log\pi_{\theta}({y_{\text{ref}}|x})}_{\text{SFT}},$ (24)
which can be seen that, no length normalization is adopted, link function $\max(0,\delta-x)$ adopts the hinge function, home advantage can be seen as None or 0, reference policy is the SFT policy to be aligned (in the regularization term), no contextual scaling, the dataset is the hybrid or offline dataset thus there is no rejection sampling. There is an SFT loss, which acts as the role of regularization.
C.3 Identity Preference Optimization (IPO)
The loss of IPO (Azar et al., 2024) is:
$\mathcal{L}_{\mathrm{IPO}}\left(\pi_{\theta};\pi_{\mathrm{ref}}\right):=\mathbb{E}_{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}\left(\beta\log\frac{\pi_{\theta}\left(y_{w}\mid x\right)}{\pi_{\text{ref }}\left(y_{w}\mid x\right)}-\beta\log\frac{\pi_{\theta}\left(y_{l}\mid x\right)}{\pi_{\text{ref }}\left(y_{l}\mid x\right)}-\frac{1}{2}\right)^{2},$ (25)
which can be seen that, no length normalization is adopted, link function $(x-1/2)^{2}$ adopts the square function, home advantage can be seen as None or 0, reference policy is the SFT policy to be aligned, no contextual scaling, the dataset is the hybrid or offline dataset thus there is no rejection sampling. Also there is no SFT loss.
C.4 CPO
CPO (Xu et al., 2024) is motivated to improve the memory and speed efficiency of DPO by neglecting the reference policy, further accompanied by a SFT loss term:
$\mathcal{L}_{\mathrm{CPO}}\left(\pi_{\theta}\right):=-\mathbb{E}_{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}\left[\log p_{\theta}(y_{w}|x)+\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}\right)\right].$ (26)
which can be seen that, no length normalization is adopted, link function $-\log\sigma$ adopts the logistic function, home advantage can be seen as None or 0, reference policy is None (or free), no contextual scaling, the dataset is the hybrid or offline dataset thus there is no rejection sampling. Also there is an SFT loss.
C.5 RSO
The loss of RSO (Rafailov et al., 2023) is:
$\mathcal{L}_{\mathrm{RSO}}\left(\pi_{\theta};\pi_{\mathrm{ref}}\right):=-\mathbb{E}_{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}_{\text{RS}}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(y_{w}\mid x\right)}{\pi_{\text{ref }}\left(y_{w}\mid x\right)}-\beta\log\frac{\pi_{\theta}\left(y_{l}\mid x\right)}{\pi_{\text{ref }}\left(y_{l}\mid x\right)}\right)\right],$ (27)
which can be seen that, it only differs from DPO in applying rejection sampling for formulating the preference dataset.
C.6 ODPO
ODPO (Amini et al., 2024) proposed to add a margin to capture the significance of preference pairs; they model this margin, or they call offset $\Delta_{r}$ as a monotonically increasing function $f(\cdot)$ of the difference between the scores associated with the responses:
$\Delta_{r}(x,y_{w},y_{l})=\alpha f\left(\operatorname{score}\left(x,y_{w}\right)-\operatorname{score}\left(x,y_{l}\right)\right),$ (28)
where $\alpha$ is a hyper-parameter that controls the extent to which an offset should be enforced. The resulting objective becomes:
$\mathcal{L}_{\mathrm{ODPO}}\left(\pi_{\theta};\pi_{\mathrm{ref}}\right):=\\ -\mathbb{E}_{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\Delta_{r}(x,y_{w},y_{l})\right)\right].$ (29)
ODPO only differs from DPO in applying a contextual dependent margin/home advantage.
C.7 ORPO
Opposed to maximizing the likelihood ratios of winning and losing answers in the preference pair in DPO, ORPO (Hong et al., 2024) propose that odds ratio can be a more sensible choice.
$\mathcal{L}_{\mathrm{ORPO}}\left(\pi_{\theta}\right):=\\ -\mathbb{E}_{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}\left[\log p_{\theta}(y_{w}|x)+\lambda\log\sigma\left(\log\frac{p_{\theta}(y_{w}|x)}{1-p_{\theta}(y_{w}|x)}-\log\frac{p_{\theta}(y_{l}|x)}{1-p_{\theta}(y_{l}|x)}\right)\right].$ (30)
where $p_{\theta}(y|x)=\exp\left(\frac{1}{|y|}\log\pi_{\theta}(y|x)\right)$. ORPO is similar to CPO in the sense that it is also reference model free and combined with a SFT loss; in addition, notably that ORPO also adopts a form of length regularization by normalizing the likelihoods with respect to the length, as in the definition of $p_{\theta}(y|x)$; finally, they compute odds ratio instead of the original likelihood ratio.
ORPO differs from DPO in being reference-model free and also yields a SFT term for regularization. Its contextual scaling is implicit, as shown in Theorem 1 for ORPO equivalent objective.
C.8 MallowsPO
Chen et al. (2024a) propose a contextual scaled objective derived from MLE under Mallows: compared to DPO that puts equal weights on each prompt and preference pairs, the resulting MallowsPO adds a contextual scaling factor $\phi(x)$ that represents this dispersion of the preferences of answers to each prompt $x$:
$\mathcal{L}_{\mathrm{MallowsPO}}\left(\pi_{\theta};\pi_{\mathrm{ref}}\right):=\\ -\mathbb{E}_{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}\left[\log\sigma\left(\phi(x)\left[\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right]\right)\right].$ (31)
To compute this dispersion, MallowsPO provided a direct approach by using a normalized predictive entropy of preference pairs $\{y_{i}^{w},y_{i}^{l}\}_{i=1,\ldots,N}$ with $N=\max(|y^{w}|,|y^{l}|)$:
$\phi(x)=-\log\left(\frac{\frac{1}{2}\sum_{i=1}^{N-1}\left[H_{\pi_{\text{ref}}}(Y_{i+1}\mid Y_{i}=y^{w}_{i})+H_{\pi_{\text{ref}}}(Y_{i+1}\mid Y_{i}=y^{l}_{i})\right]}{\log(n)}\right).$ (32)
MallowsPO differs from DPO in adding a contextual scaling factor; others are the same.
C.9 SimPO
SimPO (Meng et al., 2024) proposes a simple yet effective objective that is claimed to match or even outperform the performance of DPO:
$\mathcal{L}_{\mathrm{SimPO}}\left(\pi_{\theta}\right):=-\mathbb{E}_{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)\right],$ (33)
where $\gamma$ is introduced as a target reward margin to help separating the winning and losing responses. SimPO is similar to CPO in the sense of being reference model free; it also adopted the length normalization for the likelihoods as in ORPO; finally, it additionally introduced a constant margin to be tuned that could help to further improve the performance by encouraging a larger difference between the normalized likelihoods.
SimPO differs from DPO in adopting length normalization, a margin/home advantage term and it is also reference-model free.
Appendix D Proofs and Details
D.1 Proof of ORPO upper bound in Equation 4
Theorem 1.
$\Delta_{\theta}(x,y_{w},y_{l})=\frac{1}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{1}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)\geq 0$
$\mathcal{L}_{\textsc{ORPO-PO}}\left(\pi_{\theta}\right)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\log\frac{p_{\theta}(y_{w}|x)}{1-p_{\theta}(y_{w}|x)}-\log\frac{p_{\theta}(y_{l}|x)}{1-p_{\theta}(y_{l}|x)}\right)\right],$ (34)
has an upper bound such that
Proof.
Taking gradient of $f(x)$ with respect to $x$, we have:
Appendix E Experimental details
E.1 AlpacaEval and AlpacaEval2 Introduction
We briefly introduce AlpacaEval (Li et al., 2023) and its evaluation pipeline and metric. This part is mainly modified and summarized from descriptions in AlpacaEval official Github repository. [^5] https://github.com/tatsu-lab/alpaca_eval.
In practice, for either training a LLM or comparison between different LLMs, we need to evaluate an instruction-following model (e.g., ChatGPT). However, evaluation of such models typically requires human interactions, which is time-consuming, expensive, and hard to replicate.
AlpacaEval. AlpacaEval is an LLM-based automatic evaluation that is fast, cheap, replicable, and validated against 20K human annotations. It operates on a fixed set of 805 instructions chosen to be representative of user interactions on the Alpaca web demo, [^6] https://github.com/tatsu-lab/stanford_alpaca. and has been widely adopted for model development. Concretely, AlpacaEval adopts an automatic evaluator that has high agreement with humans (validated on 20K annotations), and evaluates a model by measuring the fraction of times a powerful LLM (e.g., GPT-4) prefers the outputs from that model over outputs from a reference model. The evaluators of AlpacaEval enable caching and output randomization by default.
AlpacaEval2. AlpacaEval2 (Dubois et al., 2024), also called Length Controlled AlpacaEval, is a length-debiased version of AlpacaEval. One of the major issues of AlpacaEval is that one can increase the win-rate by increasing the length of outputs. The main idea of Length Controlled (LC) WR is that for each model, AlpacaEval2 will fit a logistic regression to predict the preference of the autoannotator given: (1) the instruction, (2) the model, and (3) the difference of length between the baseline and model output. Given such a logistic regression AlpacaEval2 can then try to predict the counterfactual “what would the preference be if the model’s output had the same length as the baseline” by setting the length difference to 0. By averaging over this length-controlled preference, AlpacaEval2 then obtains the LC win rates.
Length controlled win-rates increase the correlation between AlpacaEval’s leaderboard and Chat Arena from 0.93 to 0.98 Spearman correlation, while significantly decreasing the length gameability of the annotator. We refer the more concrete details of the datasets, comparison of models or evaluators and leaderboards to AlpacaEval’s official website https://github.com/tatsu-lab/stanford_alpaca.
Our paper. In this paper, we adopted two evaluators: GPT4 turbo and Llama3 70B to evaluate our models’ generations against base/reference answers generated by GPT4 turbo. We adopt the default template for evaluators provided by AlpacaEval and we adopt the following fixed generation config for each model: max new tokens 4096, temperature 0.7 and top $p$ 0.1.
E.2 Training Details